The magnitude of the force developed by raising the temperature from 0 °C to 100 °C of the iron bar of 1.00 m long and 1 cm^2 cross-section when it is held so that it is not permitted to expand or bend is

(a = 10^-5/°C and Y = 1011 N/m^2)

Grade:11

2 Answers

Arun
3 years ago
The expression for force is F= Y x A x α x ΔT where Y is the Young's modulus, A the area of cross section, α the coefficient of linear expansion and ΔT is temperature difference.It is independent of length. 

F = (1011 N/m2 ) x 1 x 10−4 m2 x (10−5 /
°C) x 100 °C
   = 10
4 N
Vikas TU
3 years ago
Expression of force is F=×Y×A×α×ΔT where Y is the Young's Modulus.
∴F=10^11]cfracNm^2×1×10^−4m^2 ×(10^−5)×100°C
F=10^4 N
The magnitude of force =10^4 N